The problem should have been x^2+7x+10=0 not x^2+10=0. And the 6 was not suppose to be in the other problem.
Answer by checkley77(12569)

x^2+7x+10=0
(x+5)(x+2)=0
x+5=0
x=-5 answer.
x+2=0
x=-2 answer.
Proofs:
-5^2+7*-5+10=0
25-35+10=0
35-35=0
0=0
-2^2+7*-2+10=0
4-14+10=0
14-14=0
0=0